OpenStudy (anonymous):

Evaluate y/2 - 3 = 1 ???

OpenStudy (anonymous):

y/2=4 y=8.

OpenStudy (anonymous):

yeah its 8. y/2=4

OpenStudy (curry):

oh wati ye cause 2-3 isn't in paranthese

OpenStudy (anonymous):

y = 8

OpenStudy (anonymous):

y/2-3=1 x2 y-6=2 y=8